John takes 3 hours longer than Andrew to peel 500kg of apples.If together they can peel 500kg of apples in 8 hours,then how long would it take each one working alone?
Answer by richwmiller(17219)

1/x+1/(x+8)=1/3
(3*x*(x+8))/x+(3*x*(x+8))/(x+8)=(3*x*(x+8))/3
(3*(x+8))+3*x=x*(x+8)
3*x+24+3*x=x^2+8*x
6*x+24=x^2+8*x
x^2+2*x-24=0
Can it be factored?
What factors of -24 add up to 2?
Nevertheless, we will complete the square
x^2+2*x=24
Add it.
x^2+2*x+(2/2)=24+(2/2)
Square it.
x^2+2*x+(4/4)=24+(4/4)
x^2+2*x+(1)=24+(1)
x^2+2*x+(1)=25
(x+1)^2=25
(x+1)^2=25
We will continue.
x+1=(+-)5.0
x=-1(+-)5.0
Positive useful answer:
x=-1+5.0
One worker is x=4
x+8=8+4
x+8=12
Other worker =12
check
1/x+1/(x+8)=1/3
1/4+1/(4+8)=1/3
1/4+1/12=1/3
0.25+0.83333333e-1=0.33333333
0.33333333=0.33333333
ok
Useless answer:
x=-1-5.0
Useless answer is x=-6.0
x =-6.0
This answer is useless because it is Negative
For real life cases involving time, distance and other measurements, we reject the negative results